OpenStudy (anonymous):

y - 2x = 10 y = 4x + 8 Solve by substitution.

OpenStudy (aripotta):

well, you know that y = 4x + 8, so you just substitute that for y in the first equation

OpenStudy (anonymous):

y=2x+10 2x+10=4x+8 solve

OpenStudy (anonymous):

2x+2=4x 2x=2 x=1

OpenStudy (aripotta):

do you know what to do now?

OpenStudy (anonymous):

Actually can do do it again. I'm still lost.

OpenStudy (anonymous):

first you would find y in terms of x

OpenStudy (aripotta):

it would be (4x + 8) - 2x = 10

OpenStudy (aripotta):

so you combine like terms first

OpenStudy (aripotta):

4x and -2x being like terms

OpenStudy (anonymous):

How do you get -2x?

OpenStudy (aripotta):

because in our equation, it says (4x + 8) - 2x = 10. minus 2x is the same as -2x

OpenStudy (aripotta):

so to combine 4x and -2x, it would be 4x - 2x, which is...

OpenStudy (anonymous):

Oh man. I wrote that backwards.

OpenStudy (anonymous):

Its y = 2x - 10

OpenStudy (aripotta):

what? can you write the whole question correct? :/

OpenStudy (anonymous):

Its y = 2x - 10 y = -4x + 8

OpenStudy (aripotta):

oh

OpenStudy (aripotta):

ok, so it would be 2x - 10 = -4x + 8

OpenStudy (aripotta):

now you have to add 4x to both sides

OpenStudy (aripotta):

wait.

OpenStudy (aripotta):

nvm, continue doing that

OpenStudy (anonymous):

Alright thanks. I understand now.

OpenStudy (aripotta):

oh ok.